OpenStudy (anonymous):

The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (anonymous):

so you can get the slope of x-2y=6 with -a/b so the slop is -1/-2 or 1/2 parallel lines have the same slope so we will use this slope with point-slope form y-y1=m(x-x1) substitute the values the point is (-2,4) (x1,y1) y-4=1/2(x+2) distribute y-4=1/2x+1 add 4 to both sides y=1/2x+5 this is the line in slope-intercept form

terenzreignz (terenzreignz):

http://openstudy.com/study#/updates/52542408e4b0328cd50392ba heads up, the question has been reposted :D

terenzreignz (terenzreignz):

@austinL don't let me stop you though :D

OpenStudy (austinl):

Sorry I was on a phone call.